OpenStudy (anonymous):

Need a little help with the following questions: 1. Write an equation in standard form for a line with a slope of 1/2 and passing through A(1,2). A. x-2y-3=0 B. x-2y+3=0 C. -x+2y+3=0 D. -x-2y+3=0 2. Which is an equation for the line passing through B(0,3) and C(-3,4)? A. x-3y-9=0 B. 3x-y-3=0 C. 3x+y-3=0 D. x+3y-9=0

OpenStudy (anonymous):

I think a graph or something is needed?

OpenStudy (texaschic101):

use point slope formula y - y1 = m(x - x1) slope(m) = 1/2 (1,2) x1 = 1 and y1 = 2 now lets sub y - 2 = 1/2(x - 1) y - 2 = 1/2x - 1/2 y = 1/2x - 1/2 + 2 y = 1/2x - 1/2 + 4/2 y = 1/2x + 3/2 multiply the equation by 2 2y = x + 3 -x + 2y = 3 x - 2y = 3 x - 2y - 3 = 0 <-- answer

OpenStudy (texaschic101):

do you have any questions about this one ?

OpenStudy (anonymous):

Why do you multiply the equation by 2?

OpenStudy (texaschic101):

to get rid of the fractions you multiply the whole equation by the LCD

OpenStudy (anonymous):

Oh got it, didn't see that. Thank you so much.

OpenStudy (texaschic101):

you ready for the second problem

OpenStudy (anonymous):

I can get the slope of the two points, can I use any of the two points to sub?

OpenStudy (texaschic101):

there is a slope formula : slope(m) = (y2 - y1) / (x2 - x1) (0,3) x1 = 0 and y1 = 3 (-3,4) x2 = -3 and y2 = 4 now just sub slope(m) = (4 - 3) / (-3 - 0) can you finish and find the slope ?

OpenStudy (anonymous):

Slope would be 1/-3, if that's correct I got the rest.

OpenStudy (texaschic101):

slope is -1/3

OpenStudy (anonymous):

Alright, makes sense. Thank you

OpenStudy (texaschic101):

your very welcome....you sure you got the rest ?

OpenStudy (texaschic101):

tag me if you run into problems